Natural uranium is a mixture of 238U (99.3%, monisotopic mass = 238.05 amu) and 235U (0.7%, monisotopic mass = 235.04 amu). The 235U isotope is essential to nuclear reactors. We isolate this isotope by converting a sample that contains both forms of uranium to UF6. What is the molecular mass of the UF6 compound if only 238U is used? What is the molecular mass of the UF6 compound if only 235U is used? What would the (rate of 235U)/(rate of 238U) be according to Graham’s law?How much faster will the 235 isotope travel compared to the 238 isotope?? Show work please. I have tried many times but keeps saying my answers are incorrect.
Answer #1

atomic mass of U238 = 238.05 g

MOLECUALR MASS OF UF6 CONTAINING  U238 = 238.05 + 6 X 18.99 = 351.99 g

MOLECUALR MASS OF UF6 CONTAINING  U235 = 235.04 + 6 X 18.99 = 348.98 g

RATE OF DIFFUSION IS INVERSELY PROPORIONAL TO SUARE ROOT OF MOLAR MASS.

  (rate of 235U)/(rate of 238U) = [(MW of 238U)/(W of 235U)]1/2

= [351.99/348.98]1/2

= 1.004

ALMOST BOTH ISOTOPES TRAVEL WITH SAME SPEED.

-------------------------------------------------------------------------------------------

FOR VALUABLE INFORMATION 1 amu = 1.67377 x 10 -24 gram

to remember, 1 amu x avagadro number = 1

so that atomic masse or molar masses can be expressed in grams

for instance mola mass of U238 = 238.05 X 1.67377 x 10 -24 gram X 6.02 X 1023

= 238.05 X 1

= 238.05 g

HOPE CATCH IT.